OpenStudy (anonymous):

why can’t you use the chain rule to differentiate x/(x^2+1)?

OpenStudy (anonymous):

You can. What is chain rule btw?

OpenStudy (anonymous):

(f(x)g(x))' = f'(x)g(x) + g'(x)f(x)? is this the chain rule

OpenStudy (anonymous):

d/dx f(g(x))=f'(g(x))*g'(x)

OpenStudy (anonymous):

There is no chain there.

OpenStudy (anonymous):

f(x)/g(x)

OpenStudy (anonymous):

Oh, so that's the chain rule. Why would you use chain rule here then?

OpenStudy (anonymous):

That's a rhetorical question.

OpenStudy (anonymous):

you could techinically chain it, but you'd have to multiply too, when it's just easier to only divide

OpenStudy (apoorvk):

You can use Chain Rule on any problem anywhere in the world. (Well almost. And if the solution doesn't turn out a fifty pages long one).

OpenStudy (apoorvk):

By the way, using the quotient rule would be way easy here.

OpenStudy (anonymous):

I was thinking the inside function was u=x^2+1, and the outside function would be x/u but it has that extra x in there I don't know what to do with

OpenStudy (anonymous):

ya u can use quotient rule.............

OpenStudy (anonymous):

YOU DON'T SAY!?!

OpenStudy (anonymous):

you could use u sub if it was in an integral

OpenStudy (zarkon):

write \[x/(x^2+1)\] as \[x\cdot(x^2+1)^{-1}\] then you can use the chain rule (along with the product rule)